# Condensing Against Online Adversaries

### v1

@@ 8 Open Problems
  • Explicitly construct a seeded condenser with dependence on seed length being $1\cdot\log(1/\varepsilon)$. This will immediately make all our existential condenser constructions explicit.
@@
  • Construct a function that is $t$-online resilient for $t=\omega\left(\frac{n}{\log^{2}(n)}\right)$ or show that $O\left(\frac{n}{\log^{2}(n)}\right)$ is tight. Note that one can obtain $t=O\left(\frac{n}{\log^{2}(n)}\right)$-online resilient functions using the Ajtai-Linial function [ajtai_influence_1993].
  • Construct $\varepsilon$-collective sampling protocols with lesser number of rounds than the ones obtained using uniform oNOSF source condensers. It will also be interesting to explicitly construct such protocols when number of players is very large compared to the number of bits each player has access to. Also proving lower bounds for $\varepsilon$-collective sampling protocols will be very interesting.
Acknowledgments
We thank Rocco Servedio for extremely useful conversations about the notion of online influence. In particular, the result in Section 7.3 was obtained jointly with Rocco. We also thank Madhur Tulsiani for asking a question during our talk at FOCS’24 that motivated us to consider the model of local oNOSF sources in Section 1.5.

### v2

@@ 10 Open Problems
  • Some of our condenser results are only existential and not explicit. It would be very interesting to find explicit constructions with similar parameters. As we show, one way of achieving this would be to explicitly construct a seeded condenser with dependence on seed length being $1\cdot\log(1/\varepsilon)$.
@@
  • Construct $\varepsilon$-collective sampling protocols with fewer rounds than the ones obtained using uniform oNOSF source condensers. It will also be interesting to explicitly construct such protocols when the number of players are very large compared to the number of bits each player has access to. Further, proving lower bounds for $\varepsilon$-collective sampling protocols is a natural direction to explore.
  • Determine the exact threshold for extracting from oNOBF sources and oNOSF sources . Our lower bounds show extraction is impossible when $g\leq 0.99\ell$ while our constructions using leader election protocols require $g\geq\ell-\Omega\left(\frac{\ell}{\log\ell}\right)$ for oNOBF sources and $g\geq\ell-\Omega\left(\frac{\ell}{\log^{*}(\ell)}\right)$ for ${\left(g,\ell,n\right)}$-oNOSF sources where $n\geq\log(\ell)$. Using the connection between extractors and leader election protocols, lower bounds for extraction imply lower bounds for leader election protocols. In particular, matching lower bounds for extraction would imply all current leader election protocols are tight, a long standing open problem.
Acknowledgements
We thank Madhur Tulsiani for asking a question that motivated us to consider the model of local oNOSF sources in Appendix A. We thank the organizers of the Dagstuhl Seminar on Algebraic and Analytic Methods in Computational Complexity and Schloss Dagstuhl for providing a stimulating research environment, where discussions between R.S. and E.C. contributed to this collaboration.
@@ Acknowledgements
Appendix A Extracting from Local oNOSF Sources
Definition A.1 (Local oNOSF sources ).
[…]
Theorem A.2 (Using the explicit extractor of [chattopadhyay_extractors_2022]).